# Evaluation of ChatGPT Usability as A Code Generation Tool

### v2

@@ III Methodology
In this section, we propose a user centric approach to address the above problems.
@@
The general principle of the approach is that, first, the test dataset should be constructed according to the use scenario. Second, the process of testing the ML model on each test case should be resemble how the users use it. Moreover, the outputs from the ML model should be evaluated on quality attributes that reflect the usability of the output from the users’ point of view. Finally, the performance metrics should represent user’s experience with using the ML model. By doing so, we put the user at the center of the evaluation, thus the name of user centric evaluation.
We will first outline the key ideas of the method in subsection III-A. We will then delineate the method with the evaluation of ChatGPT on its usability of generating R program code.
III-A Main Ideas
Our new methodology consists of the following new features.
III-A1 Multi-attempt testing process
@@
In the testing of a LLM, we require the testers to try their best to complete the specified task with as few attempts as possible. Each attempt consists of formulating the best text inputs to invoke the LLM and then inspecting the solution generated by the LLM. If the output from the LLM is unsatisfactory, a second attempt is made. The process continues until either the number of attempts reached the maximum of $k$, or a satisfactory solution is obtained. This leads to two performance metrics: average number of attempts, denoted by #$\#attempt_{k}$, and average time used to complete the test tasks, or average completion time for short. They reflect the user experiences in the use of the LLM. The testing process better matches the process that a user typically uses a LLM to generate code than the single attempt approach and the corresponding performance metric $pass@k$.
@@
It is worth noting that being user centric, whether a solution is satisfactory should be judged by the user. In our experiment, a guideline is given to the tester that a solution should be judged as satisfactory if it is correct when compared with the standard reference solution, or it is very close to the reference solution. A functionally incorrect solution may well be regarded as satisfactory if a minor editing of the machine generated solution will work. For this reason, when the number of attempts reached the maximum threshold, the test should not be considered as a complete failure. Instead, the quality of the result at the final attempt should also be assessed.
III-A2 Metadata to describe test case’s usage
@@
Our methodology to the evaluation of ChatGPT’s usability as a code generation tool is a scenario-based testing approach. Conceptually, scenario-based testing is to test a software on a series of subsets of test sets that each subset represents a specific scenario of using the system. Evaluating a ML model’s performance on each subset separately can provide the detailed performance information for different usages of the system. As reported in [36], this will also enable us to identify the strength and weakness of the model in different usage scenarios, thus providing directions for future improvement of the model. However, scenario combinations may cause duplicated running of test cases. To support efficient scenario-based testing and evaluating a ML model, we propose to use metadata associated with each data item in the dataset to describe the usage scenarios that the data item represents. This can eliminate the need for duplicated execution of test cases. This paper will demonstrate this by combining the scenarios that the task is of different difficult levels with scenarios to generate different types of codes.
III-A3 A set of quality attributes on usability
Based on the assumption that the code generated by a LLM will be examined and possibly revised by the human user, we propose a set of quality attributes on the usability of the solution generated by the LLM. These quality attributes are not only on the generated program code but also on the generated texts that explain the code. They reflect various aspects of how close the output of LLM is to the perfect solution, how easy to revise it to a usable solution, and how good the output is in the context of code maintenance, debugging, evolution and reuse. These attributes include accuracy, readability, structuredness, logic clarity, depth of explanations, completeness, and coverage of parameters, etc. Although such attributes heavily rely on the tester’s subject assessment, they better reflect users’ true opinions on the usability of the tool. To complement such subjective metrics, we also measure the time taken to complete the task and the number of attempts the tester made to finish the task. These are objective measurements.
III-A4 Multi-scored metrics
For each quality metric, the output from the LLM is assessed on the scale from 1 to 5, where 1 is the poorest and 5 the best. Each score value between 1 and 5 is clearly defined for each quality attribute to minimize the impact of subjectiveness in the assessment. It also reflects the fact that, even if a generated code is imperfect, it could still be useful after manual revision.
III-A5 A sequential multi-stage process to evaluate learnability
To study the learnability of LLMs, each tester engages in a linear sequence of experiments on disjoint subsets of the test dataset, which are drawn at random from a given dataset. The performance of the system is evaluated statistically on each subset separately and compared to understand the trend of performance improvement. This enables us to study the learnability of the LLM.
@@ III-A5 A sequential multi-stage process to evaluate learnability
In the remainder of the section, we delineate the proposed method with the evaluation of ChatGPT on its usability of generating R program code. We will present the research question addressed, the design of the experiment, the construction of our benchmark dataset and the performance evaluation criteria applied. The results of the experiment will be presented in Section IV.
III-B Research Questions and Experiment Design
@@
Our goal is to evaluate ChatGPT’s usability as a tool of generating program code in R language. The following are the specific research questions.
@@
  • RQ0: How good is the usability of ChatGPT as a R program code generation tool? Here, usability is measured in terms of (a) the quality of the generated code as measured on various usability attributes, (b) the user’s experiences in using ChatGPT as measured by the average number of attempts, and the average completion time.
@@
  • RQ3: How easy to learn the uses of ChatGPT as a code generation tool?
@@
To answer these research questions, we first constructed a benchmark. Then, five disjoint test subsets $T_{0},T_{1},\cdots,T_{4}$ were drawn at random from the benchmark. These test sets are of equal size, which consists of 20 test cases. Since these test sets are drawn at random, each of them contains various types of questions and various levels of difficulty, and from different sources.
@@
The experiments are conducted with the same subject (i.e., a human tester) who is a master degree student studying on a data analytics program but has no experiences in the use of ChatGPT before the experiment. The testing of ChatGPT is performed sequentially from $T_{0}$ to $T_{4}$.
@@
In the experiment, each test case is a task to be completed by the tester by using ChatGPT. The tester is required to formulate a query to invoke ChatGPT with the coding question. The output from ChatGPT is recorded and checked. If the output from ChatGPT is not satisfactory in terms of the correctness of the generated code, a further prompt is submitted to ChatGPT. This is repeated until a satisfactory output is obtained or it is terminated after 10 cycles of input/output. Each cycle is called an attempt. The time used to complete the task on each test case and the number of cycles engaged with ChatGPT were also recorded. The final output is then assessed on a set of quality criteria presented in the next subsection.
@@
Statistical analysis is conducted on the union $T$ of these test sets, i.e., $T=T_{0}\cup T_{1}\cup\cdots\cup T_{4}$, for the overall performance of ChatGPT to answer research question RQ0. The test set $T$ is then split into subsets according to difficulty level and types of program code to answer research questions RQ1 and RQ2 using the metadata associated with the test cases without re-executing them. To answer research question RQ3, the tendency of performances changes is analyzed separately on the disjoint subsets $T_{0},T_{1},\cdots,T_{4}$.
III-C Construction of Dataset
@@
As far as we know, there is no existing dataset available for generating R program codes. Being a special purpose language targeting data analysis and statistics, R contains many special features. General datasets of code generation and for other languages cannot provide a fair evaluation. Therefore, we constructed a dataset specifically for the generation of R program code.
III-C1 Source of Data
@@
The data are gathered from exercise questions of accredited R programming textbooks, which covers a wide range of topics, including uses of various language facilities and libraries, numerical calculations, data structural manipulations and transformations, data visualization and statistical analysis, etc. The questions are of different difficulty levels. The questions are well presented as they are from well written textbooks. These books are primarily used in the IT industry to educate and train professional programmers who have no experience of programming in the R language. We have collected all questions from six textbooks given in references [37, 38, 39, 40, 41, 42]. They are referred to as B1 to B6, respectively, in the sequel.
@@ III-C2 Metadata
  • Easy: when the program code to be generated consists of a small number of lines of code that is essentially a simple invocation of a library function.
  • Medium: when the program code to be generated has about 10 lines of code that is beyond a simple invocation of predefined functions. However, it does not involve complicated data structure and algorithms.
  • Hard: when the program code to be generated is more complicated than ten lines and requires more complex data structures and algorithms.
@@
(c) Sources of the data. The source of a question is also provided as a metadata of the test case. Since our questions are all from textbooks, the metadata contains the title of the book, the year of publication and the chapter of the book where the original question is in. This metadata is potentially useful to analyze how well a LLM deals with new programming tasks and uses new libraries and facilities of the R programming language, although it is not used in the experiments reported in this paper.
The data are represented in JSON format. Figure 1 shows the structure of the JSON representation of test cases. An example is given in Figure 2.
@@
The dataset contains 351 test cases in total. Table 1 shows the statistical characteristics of the dataset. Each row of the table gives the number of test cases of a specific feature given in the first column from different sources in column B1 to B6 and the total number in the column Total. The final row shows the total number of test cases from each source.
TABLE II: Statistical Characteristics of the Dataset
@@
III-D Quality Attributes and Evaluation Criteria
The solution produced by ChatGPT is assessed on the following criteria.
(1) Accuracy: The correctness and accuracy of the LLM model’s response in comparison with the standard answer to the programming question. It is assessed through exe-cuting the code produced by ChatGPT and the standard answer to the coding question. Their outputs are compared to see if the generated code gives correct output.
(2) Completeness: the completeness of the solution that covers the requirements of the programming question. It is assessed by comparing the functionality of the code gen-erated against the coding problem.
(3) Structuredness: the structure of the code generated and formatted. It is assessed by inspection of the code on how well the code is structured and organized.
(4) Conciseness: The conciseness of the code produced by ChatGPT without compromising functionality. This is assessed by review of the code in comparison with the standard answer. Typically, the generated code and standard answer are compared on the number of lines of code.
(5) Logic clarity: the coherence, soundness, and clarity of the generated code. This is examined and assessed via review of the code to identify inconsistencies or logic errors.
(6) Parameter Coverage: how well the parameters pro-vided in the coding question are used or covered in the solution produced by the LLM model. Here, “parameter” refers to the coding elements given in the question like dataset name, variable name, data values, function names, library code names, etc. This is assessed by inspection of the code to measure if all coding elements in the question were used and appropriately incorporated in the solution.
(7) Readability: the easiness for human readers to read and understand the produced solution. This is assessed by a human tester to read the produced solution and attempt to understand it, where a solution typically includes pieces of code together with natural language explanations.
(8) Depth of Explanation: the quality of the textual explanation produced together with the code. This is assessed by reviewing the text produced by the language model accompanying the code and provides explanation of the generated code. It focuses on how well the text explains the code in a comprehensible and understandable way and to elucidate the rationale behind the code.
For each criterion, the performance of the LLM model is scored in the scale from 1 to 5, where 1 is the worst and 5 is the best. Table III gives the criteria to assess the logic clarity of the generated solution. The criteria for other quality attributes can be found in Appendix A.
TABLE III: Assessment Criteria for Logic Clarity
Score Criterion
1
The response lacks logic and has no clear flow of thought as demonstrated in the code.
2
The response demonstrates some flow of thought, but it contains many lines of code misplaced.
3
The response demonstrates a flow of thought, but some lines of code are misplaced.
4
The response demonstrates a clear logic flow of thought but could be improved by minor rearrangement of the code.
5
The response demonstrates a clear flow of thought, and the code is logically well ordered.
IV Experiment Results
This section reports the experiment results.
IV-A Overall Performance
To answer research question RQ0, the overall performance of ChatGPT on its capability of generating R language code is analysed on all test cases in the test set $T=T_{0}\cup\cdots\cup T_{4}$. The experiment results are shown in Figure 3 with the average performance scores on various quality criteria.
@@ IV-A Overall Performance
The time to complete the task is mostly spent on formulating text inputs as the queries to the ChatGPT and reviewing the output obtained from ChatGPT. ChatGPT responds to queries so fast that the time is negligible. Figure 5 shows the distribution of the time needed to complete the task of generating R code, where X-axis is the time required to complete the task of generating R code, Y-axis is the number of test cases. The average of completion time is only 47 seconds while the median is 31.50 seconds. It is worth noting that 90% of test cases can be completed within 100 seconds, while 98% of test cases were completed within two and half minutes. The two cases that required more than 5 attempts took more than 180 seconds, where one of them was given up after 10 attempts, which took 462 seconds, which is less than 8 minutes.
@@
The objective metrics of performance confirmed our conclusion that ChatGPT’s capability of code generation is satisfactory in general.
IV-B Types of Code to Generate
A detailed analysis of ChatGPT’s performance on generating different types of R code reveals that it has the best performance on generating general programming type of code, while weakest in generating visualization and exploration type of code. The average total performance scores on different types of questions are depicted in Figure 6.
Figure 6: Average Total Scores of Various Types of Questions.
The data shows that ChatGPT scored an average of 38.08 out of 40 (i.e. 95.2%) on general programming questions, 36.43 (91.1%) on visualization and 36.63 (91.6%) on exploration.
@@ IV-B Types of Code to Generate
Figure 7: Average Numbers of Attempts to Complete Tasks of Generating Various Types of Code.
Figure 8: Average Completion Times on Tasks of Generating Various Types of Code (Seconds).
@@
Figure 9: Distributions of The Numbers (%) of Attempts to Complete Tasks of Generating Various Types of Code.
Figure 10: Distribution of Task Completion Times on Tasks of Generating Various Types of Code.
Figure 11 shows the average scores of various quality attributes for generating visualization type of code. The data shows that conciseness is the weakest aspect of quality on generating visualization code, which scored 3.43 out of 5, i.e. 66.8%.
Figure 11: Average Performance Scores on Visualization Questions in Various Quality Attributes.
IV-C Difficulty Levels
Will ChatGPT’s usability decrease as the difficulty of the task increases? Yes. The experiment data shows clearly that the total quality scores on subjective quality criteria decrease as the level of difficulty increases, while the average number of attempts and the average amount of time required to complete code generation tasks increase as the difficulty level increases; see Figure 12.

### v3

@@ III The Proposed Methodology
To address the above problems, we propose a user centric approach to the evaluation of LLMs as intelligent coding tools.
@@
The general principle of the approach is that, first, the test dataset should be constructed according to the use scenario in which the evaluation is performed. Second, the process of testing and evaluation on each test case should resemble the process in which the ML model is used as a tool. Moreover, the outputs from the ML model should be evaluated on quality attributes that reflect the usability of the output from the users’ point of view. Finally, the performance metrics should represent the user’s experience with using the ML model. By doing so, we put the user at the center of the evaluation, thus the name of user centric evaluation.
In this section, we will outline our methodology that realises these principles. It will then be delineated and explained in the case study with the evaluation of ChatGPT on its usability of generating R program code.
III-A Multi-attempt Testing Process
@@
In the testing of a LLM, we require the testers to try their best to complete the specified task with as few attempts as possible. Each attempt consists of formulating the best text inputs to invoke the LLM and then inspecting the solution generated by the LLM. If the output from the LLM is unsatisfactory, a second attempt is made by amending the input. The process continues until either the number of attempts reaches the maximum of $k$ of allowed attempts, or a satisfactory solution is obtained. This leads to two performance metrics: number of attempts, denoted by #$\#attempt_{k}$, and time used to complete the test tasks, or completion time for short. They reflect the user experiences in the use of the LLM. This testing process better matches the process that a user typically uses a LLM to generate code than the single-attempt / multi-trial approach and the corresponding performance metric $pass@k$.
@@
It is worth noting that being user centric, whether a solution is satisfactory should be judged by the user. In our case study reported in Section IV and V, a guideline is given to the tester that a solution should be judged as satisfactory if it is correct, or very close to the reference solution. A functionally incorrect solution may well be regarded as satisfactory if a minor editing of the machine generated solution will work. For this reason, when the number of attempts reached the maximum threshold, the test should not be considered as a complete failure. Instead, the quality of the result at the final attempt was also assessed in the case study. On the other hand, even if a generated solution is functionally correct, it may not be regarded as satisfactory, because to be useful it should also meet other quality requirements. which will be discussed in subsection III-C.
III-B Benchmark with Metadata
@@
For an evaluation of a LLM to faithfully reflect its users’ usage, the test dataset must represent user’s uses of the LLM. Our methodology achieves this purpose by constructing a benchmark that supports scenario-based testing in which the LLMs are tested on a series of subsets of the benchmark that each subset represents a specific scenario of using the system. Evaluating a ML model’s performance on each subset separately can provide the detailed performance information for different usages of the system. As reported in [36], this will also enable us to identify the strength and weakness of the model in different usage scenarios, thus providing directions for future improvement of the model. However, scenario combinations may cause duplicated running of test cases. To support efficient scenario-based testing and evaluating a ML model, we propose to associate metadata to each test case in the benchmark to describe the usage scenarios that the test case represents. This can eliminate the need for duplicated execution of test cases and form subsets for different scenarios easily. This is demonstrated in our case study by combining the scenarios of different difficult levels with the scenarios of generating different types of code.
III-C Quality Attributes for Assessing Usability
In the real use of a LLM for code generation, a solution produced by a LLM will be examined to determine if it is satisfactory and possibly revised by the human user after accepting the suggested solution. Therefore, the evaluation of a LLM should not only assess the correctness of the generated code, but also its usability. Here, we propose to define a set of quality attributes for this purpose. These quality attributes should be not only on the generated program code but also on the generated texts that explain the code. They should reflect various aspects of examination and revision of the generated solution, such as how easy to understand the generated solution, how easy to revise into a usable solution, and how good the output is in the context of code maintenance, debugging, evolution and reuse.
For example, the usability attributes used in our case study of ChatGPT on its capability of generating R program code is summarised in Table II. For each of these quality attributes, the output from the LLM is assessed on the scale from 1 to 5, where 1 is the poorest and 5 the best. Although such quality attributes heavily rely on the tester’s subjective assessment, they better reflect users’ true opinions on the usability of the tool.
TABLE II: Attributes of Usability Used in Case Study
Attribute Meanings
Accuracy How close the generated code to be correct
Completeness How complete the generated solution solves the programming problem
Conciseness How concise the generated solution is
Logic clarity How clear the code and explanation are presented with a clear and logical flow of thought
Readability How easy to read and understand the generated solution
Structuredness How well the code is structured
Coverage of parameters How well the parameters provided in the coding problem are used and covered in the solution
Depth of explanations How thorough the code is explained by the text in the solution
IV Case Study
@@
In this section, we delineate the proposed method by a case study with the evaluation of ChatGPT on its usability as a code generation tool for programming in the R language. We will present the research questions addressed, the design of the experiments, and the construction of a benchmark dataset. The results will be presented in Section V.
IV-A Research Questions
@@
Our goal of the experiment is to evaluate ChatGPT’s usability as a tool of generating program code in R language. The following are the specific research questions.
@@
  • RQ0: How good is the usability of ChatGPT as a R program code generation tool? Here, usability is measured in terms of (a) the quality of the generated code as measured on various usability attributes, (b) the user’s experiences in using ChatGPT as measured by the average number of attempts, and the average task completion time.
@@
IV-B Design of the Experiment
@@
To answer these research questions, we first constructed a benchmark by extracting data from various textbooks on R programming and associated metadata about the source, type of code to be generated, and level of difficulty to each test case in the benchmark. Then, five disjoint test subsets $T_{0},T_{1},\cdots,T_{4}$ were drawn at random from the benchmark. These test datasets are of equal size that each consists of 20 test cases. Since these test sets are drawn at random, each of them contains various types of questions and various levels of difficulty, and from different sources.
@@
The experiments are conducted with the same subject (i.e., a human tester) who is a master’s degree student studying on a data analytics program but has no experiences in the use of ChatGPT before the experiment. The testing of ChatGPT is performed sequentially from $T_{0}$ to $T_{4}$.
@@
In the experiment, each test case is a programming task to be completed by the tester via using ChatGPT. The tester is required to formulate a query to invoke ChatGPT with the coding question. The output from ChatGPT is recorded and checked. If the output from ChatGPT is not satisfactory in terms of the correctness of the generated code, a further prompt is submitted to ChatGPT. This is repeated until a satisfactory output is obtained or it is terminated after 10 cycles of input/output. Each cycle is called an attempt. The time used to complete the task on each test case and the number of attempts engaged with ChatGPT were also recorded. The final output is then manually assessed on the quality attributes presented in Table II.
@@
Statistical analysis is conducted on the union $T$ of these test sets, i.e., $T=T_{0}\cup T_{1}\cup\cdots\cup T_{4}$, for the overall performance of ChatGPT to answer research question RQ0. The test set $T$ is then split into subsets according to difficulty levels and types of program codes to answer research questions RQ1 and RQ2 using the metadata associated with the test cases without re-executing them and re-assessment of quality. To answer research question RQ3, the trends of performances changes are analysed separately on the disjoint subsets $T_{0},T_{1},\cdots,T_{4}$.
IV-C Construction of Dataset
@@
As far as we know, there is no existing dataset available for generating R program codes. Being a special purpose language targeting data analysis and statistics, R contains many special features. General datasets of code generation and those for other languages cannot provide a fair evaluation. Therefore, we constructed a dataset specifically for the generation of R program code.
IV-C1 Source of Data
@@
The data are gathered from exercise questions of accredited R programming textbooks, which covers a wide range of topics, including uses of various language facilities and libraries, numerical calculations, data structural manipulations and transformations, data visualization and statistical analysis, etc. The questions are of different difficulty levels. These questions are well presented since they are from well written textbooks. These books are primarily used in the IT industry to educate and train professional programmers who have no experience of programming in the R language. We have collected all questions from six textbooks given in references [37, 38, 39, 40, 41, 42]. They are referred to as B1 to B6, respectively, in the sequel.
@@ IV-C2 Metadata
  • Easy: when the program code to be generated is essentially a simple invocation of a library function.
  • Medium: when the program code to be generated is beyond a simple invocation of predefined functions. However, it does not involve complicated data structure and algorithms.
  • Hard: when the program code to be generated requires complex data structures and algorithms.
@@
(c) Sources of the data. The source of a question is also provided as a metadata of the test case. Since our questions are all from textbooks, the metadata contains the title of the book, the year of publication and the chapter of the book where the original question is in. This metadata is potentially useful to analyse how well a LLM deals with new programming tasks and uses new libraries and facilities of the R programming language, although it is not used in the experiments reported in this paper.
The metadata are represented in JSON format. Figure 1 shows the structure of the JSON representation of test cases. An example is given in Figure 2.
@@
The dataset contains 351 test cases in total. Table 1 shows the statistical characteristics of the dataset. Each row of the table gives the number of test cases in the benchmark, where the first column is the feature, columns B1 to B6 are the numbers from different sources, and the total number in the column Total. The final row of the table shows the total number of test cases from each source.
TABLE III: Statistical Characteristics of the Dataset
@@
V Experiment Results
This section reports the experiment results on the overall performance of ChatGPT on generating R program code, and the performances in various scenarios in which different types of code are to be generated, and the impacts of difficulty level on performance. Finally, the potential threats to the validity and generalisability of the experiment results will be discussed.
V-A Overall Performance
To answer research question RQ0, the overall performance of ChatGPT on its usability as a R code generator is analysed on all test cases in the test set $T=T_{0}\cup\cdots\cup T_{4}$. The experiment results are shown in Figure 3 with the average performance scores on various quality criteria.
@@ V-A Overall Performance
The time to complete a task is mostly spent on formulating text inputs as the queries to the ChatGPT and reviewing the output obtained from ChatGPT. ChatGPT responds to queries so fast that the time is negligible. Figure 5 shows the distribution of the time needed to complete the task of generating R code, where X-axis is the time required to complete the task, Y-axis is the number of test cases. The average of completion time is only 47 seconds while the median is 31.50 seconds. It is worth noting that 90% of test cases can be completed within 100 seconds, while 98% of test cases were completed within two and half minutes. The two cases that required more than 5 attempts took more than 180 seconds, where the one that was given up after 10 attempts took 462 seconds.
@@
The objective metrics of performance confirmed our conclusion that ChatGPT’s usability as a code generation tool is satisfactory in general.
V-B Performance of Generating Different Types of Code
A detailed analysis of ChatGPT’s performance on generating different types of R code reveals that it has the best performance on generating general programming type of code with an average total quality score of 38.08 out of 40 (i.e. 95.2%), while weakest in generating visualization and exploration types of code with the average total quality scores of 36.43 (91.1%) and 36.63 (91.6%), respectively. The results are depicted in Figure 6.
Figure 6: Average Total Quality Scores of Various Types of Questions
@@ V-B Performance of Generating Different Types of Code
Figure 7: Average Numbers of Attempts for Different Types of Code
Figure 8: Average Completion Times for Different Types of Code
@@
Figure 9: Distributions of The Numbers ofAttempts for Different Types of Code
Figure 10: Distributions of Task Completion Times for Different Types of Code
Figure 11 shows the average quality scores for visualization type of code on various quality attributes. The data shows that conciseness is the weakest aspect of quality on generating visualization code, which scored 3.43 out of 5, i.e. 66.8%.
Figure 11: Average Quality Scores on Visualization Questions
V-C Impacts of Difficulty Levels on Performance
Will ChatGPT’s usability decrease as the difficulty of the task increases? The experiment data shows clearly that the average total quality scores on subjective quality criteria decrease as the level of difficulty increases, while the average number of attempts and the average task complete time increase as the difficulty level increases; see Figure 12.